At a school's open house, T-shirts and sweatshirts were sold.
Each item was purchased by either a student or a parent.
The two-way frequency table summarizes a random sample of

80

items sold that night.

Student Parent
T-shirt

24

19

Sweatshirt

16

21

Let sweatshirt be the event that an item randomly chosen from among those sold was a sweatshirt.

Let student be the event that an item randomly chosen from among those sold was purchased by a student.

Find the following probabilities. Write your answers as decimals.


(a)

=P(student)

(b)

P(sweatshirtand=student)

(c)

P(sweatshirt|=student)
Answer #1
Student Parent Total
T-shirt 24 19 43
Sweatshirt 16 21 37
Total 40 40 80

a. P(Student)=\frac{40}{80}=0.5

b. P(sweat shirt and Student)=\frac{16}{80}=0.2

c. P(\frac{sweat shirt}{student})=\frac{P(Sweat shirt and Student)}{P(Student)}=\frac{0.2}{0.5}=0.4